Plant scientists developed different varieties of corns that have a rich content of lysine which is a nutritious animal feed. A group of chicks were given this food to test the quality. The distribution of the weight gains (in grams) of these chicks are shown below:


Weight gains (in grams) Frequency

318 - 335 4

336 - 353 5

354 - 371 2

372 - 389 3

390 – 407 2

408 – 425 3

426 - 443 1


Find:

(a) the mean weight gains

(b) the median

(c) the variance for the above frequency intervals

(d) the standard deviation


1
Expert's answer
2021-04-14T13:59:56-0400

We have that




(a) the mean:


"\bar x =\frac{\sum fx}{n}=\frac{7376}{20}=368.8"

(b) the median:

"\frac{n}{2}=\frac{20}{2}=10" thus class median is the 3rd class


"median=L_m+i(\frac{\frac{n}{2}-F}{f_m})"


where

n = the total frequency = 20

F = the cumulative frequency before class median = 9

i = the class width = 8

fm = the frequency of the class median = 2

Lm= the lower boundary of the class median = 353.5

"median=353.5+8\cdot(\frac{10-9}{2})=357.5"

(c) the variance for sample data is calculated by the formula:


"s^2=\frac{\sum fx^2-\frac{(\sum fx)^2}{n}}{n-1}=\frac{2743775-\frac{54405376}{20}}{19}=1237.2"

(d) the standard deviation:

"s=\sqrt {s^2}=\sqrt{1237.2}=35.2"


Answer:

(a) the mean = 368.8

(b) the median = 357.5

(c) the variance for the above frequency intervals = 1237.2

(d) the standard deviation = 35.2